if sqrt(x+y) + sqrt(y-x)=c, d2y/dx2=?
 
can anyone  plzz solve this step by step..plzzzz

Grade:12

1 Answers

sonika p
6 years ago
sqrt(x+y) +sqrt(y-x) = c
squaring on both sides, 
x+y + y-x + 2* sqrt((x+y)*(y-x)) = c2
x+y + y-x + 2* sqrt(y2  – x2) =  c2
2*sqrt(y2  – x2) =  c2 – 2y
squaring on both sides, 
4(y2  – x2) = c4 + 4y2 – 4c2y
-4x2 = c4  – 4c2y
y = (c2/4)+ (x2)/c2
dy/dx = (2x)/c2
d2y/dx2 = 2/c2